OpenStudy (anonymous):

What are the possible number of positive, negative, and complex zeros of f(x) = 2x^3 – 5x^2 + 6x – 4 ?

OpenStudy (anonymous):

[x = (1/6)*(71+6*sqrt(177))^(1/3)-11/(6*(71+6*sqrt(177))^(1/3))+5/6], [x = -(1/12)*(71+6*sqrt(177))^(1/3)+11/(12*(71+6*sqrt(177))^(1/3))+5/6+(1/2*I)*sqrt(3)*((1/6)*(71+6*sqrt(177))^(1/3)+11/(6*(71+6*sqrt(177))^(1/3)))], [x = -(1/12)*(71+6*sqrt(177))^(1/3)+11/(12*(71+6*sqrt(177))^(1/3))+5/6-(1/2*I)*sqrt(3)*((1/6)*(71+6*sqrt(177))^(1/3)+11/(6*(71+6*sqrt(177))^(1/3)))]